What is the magnitude and direction of the electric field at .35 m from a -6.5 x 10^-8 C point charge?

1
Expert's answer
2022-02-28T10:42:53-0500

The electric field due point charge

"E=k\frac{q}{r^2}\\
=\rm9*10^9*\frac{6.5*10^{-8}}{0.25^2}=9360\: N/C"

Field directed radially toward the charge.